OpenStudy (anonymous):

c^2 – 4c = 0 How do I solve for c?

OpenStudy (shaik0124):

c*c-4c+4c=0+4c c*c=4c divide by c on both sides now u will get c=4

OpenStudy (anonymous):

thanks :)